shows two parallel nonconducting rings with their central axes alono a common line. Ring i has uniform charge o, and radus R: ring 2 has uniform charge oz and the same radius R. The rings are separated by a distance 3.00R. The ratio of the electric field magnitudes of Ring 1 and Ring 2 at point P on the common line is 1.23. What is the ratio of charge maonitudes as/92? Ring I Ring 2 Number Units the tolerance is +/-5%
Answer #1

Every point on ring 1 is at distance √(R^2 + R^2) = R√2 away from P.
So if we observe a little bit of charge "q", then it produces field of magnitude
E = kq / (R√2)^2 = kq / (2R^2)
at point P

There is no reason for the resulting field to point in any direction off the line joining the two rings.
So the components of field perpendicular to line joining them cancel out. We're only interested in the horizontal components.
Each little bit of charge on ring 1 produces a field E that points at 45° from the horizontal.
The purely horizontal component of that field is E*cos45° or using similar triangles E * (√2 / 2)
= kq/R^2 * (√2 / 4)

Since it is the same for every point on the ring, add that up around the ring and you get magnitude of field produced by the entire ring.
E1 = k(q1)/R^2 * (√2 / 4)



Repeat same for ring 2.
Field produced by a little bit of charge is E = kq / ( √[R^2 + (2R)^2] )^2
E = kq / (5R^2)

Using similar triangles, horizontal component = E * (2R) / √[R^2 + (2R)^2]
= E * 2 / √5
= kq / (5R^2) * (2 / √5)
= kq/R^2 * 2/(5√5)

So magnitude of total field produced by ring 2 is
E2 = k(q2)/R^2 * 2/(5√5)



We know
E1 = 1.23 E2
Substitute and simplify
q1 * √2 / 4 = 1.23 q2 * 2/(5√5)
q1/q2 = 1.23 * 8 / (5√10) = 6.223